OpenStudy (anonymous):

how to Evaluate the function at the indicated values?? f(x)= 2x-1 a.f(x+1) b.f(2x) c.2f(x) d.f(x+h) e.f(x)+f(h)

OpenStudy (campbell_st):

just substitute f(x+1) = 2(x + 1) -1 distribute and simplify f(2x) = 2(2x) -1 distribute and simplify hope this helps you to get the idea of what you need to do.

OpenStudy (anonymous):

i'll just substitute all upto e??

OpenStudy (anonymous):

and i will get the answer?

OpenStudy (campbell_st):

yep... but you will need to distribute where necessary and simplify where necessary and (c)is 2f(x) = 2(2x -1)

OpenStudy (anonymous):

uhh ok i'll try it :) thanks

OpenStudy (cggurumanjunath):